OpenStudy (anonymous):

What is the slope-intercept form equation of the line that passes through (2, 4) and (4, 10)?

OpenStudy (anonymous):

A. y = −3x − 2 B. y = −3x + 2 C. y = 3x + 2 D. y = 3x − 2

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

first we find the slope by using slope formula slope (m) = (y2 - y1) / (x2 - x1) (2,4)...x1 = 2 and y1 = 4 (4,10)...x2 = 4 and y2 = 10 mow we sub slope(m) = (10 - 4) / (4 - 2) slope(m) = 6/2 slope(m) = 3 now we use y = mx + b we already know are slope(m) = 3 and we can use either set of points for out x and y I will use (2,4)...x= 2 and y = 4 so all we need to find is b, our y intercept now we sub 4 = 3(2) + b 4 = 6 + b 4 - 6 = b -2 = b so our equation is : y = 3x - 2

OpenStudy (anonymous):

Thank you so much! You are the best! I wish I could reward you with like money or something haha!

OpenStudy (texaschic101):

lol....thank you is good enough reward for me :)

OpenStudy (anonymous):

You're so kind. :D

OpenStudy (texaschic101):

I try :)